OpenStudy (anonymous):

In a 45°- 45°- 90° right triangle, one of the legs is 6 inches. How long are the other two sides? i know for sure it will be 6 and a √2 unsure if there should be a 6 outside the sqrt

OpenStudy (accessdenied):

 You could always try using Pythagorean theorem to confirm.

OpenStudy (anonymous):

6 and \[ 6 \sqrt 2 \]

OpenStudy (anonymous):

yeah pythagorean thank you

OpenStudy (accessdenied):

You're welcome. :) Yeah, the general case can be found using Pythagorean as well: \( h^2 = l^2 + l^2 \\ = 2l^2 \) \(h = \sqrt{2l^2}\\ = l\sqrt{2} \)